SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (, 2007)

Attendees: Debbie Bard, Rainer Bartoldus, Brandon Drummond, Al Eisner, Fang Fang, Kim Hojeong, Arik Kreisel, Rahmat Rahmat, Mark Tibbetts.

Operations

  • No beams the last days as PEP work to fix a vacuum leak in the Q4 magnets (see here and here for the latest news).
  • We also had downtime last week because of a water leak near where the pipes were replaced in the last 3 day ROD. The GLT transmission errors that we had been seeing disapeared when we started back up again, which is good but puzzling. It's possible PEP just had particularly bad trickle injection during that week of running that caused the errors, Debbie is still looking into the cause.
  • Rainer was asked for a trigger mask that would reduce the logging rate while running over the weekend taking COSMICS_NODCH. Jim also contacted us to ask if the FastMon error messages complaining about no DCH triggers could be removed for this configuration. It is hoped that we will not run in COSMICS_NODCH for much longer (it was never meant to be used as a long-term configuration).

EMT

  • TPBs have been looked over by Lupe. New chips have been ordered but the pads were knocked off with the chips so it's more complicated (but possible) to re-attatch them. TPBs can be tested at the test stand in the future.

L3

  • [Folder of plots] Arik is giving a talk at the Wednesday physics meeting this week. Essentially, the new L3 tracking makes a great improvement in the reco tracks before TrkFixUp is applied (see plot) but makes little difference after it is applied (see plot). Reco processing is faster, but efficiency is not improved (nor is it worsened!). Looking at the track-finder history you can see the improvement in processing is because more tracks are found with the L3 track converter, rather than with the backup methods. Also interesting to look at the change in hit types - tracks move from "unused" to "inactive" with the new L3 tracking.
    Future plans: Do a study on filters, decide whether to include this as part of any reprocessing at the reco level (probably not).

Trickle injection

  • Matt Weaver has produced a module to monitor trickle injection quality for MCC.

AOB

  • Lumi - see Grant's posting about non-statistical variations in backgrounds in mu-mu MC.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.